The plasma density in the lower ionosphere has
been measured during satellite re-entry to be
about 1018 m-3 at 50km altitude, 1017 at 70 km,
and 1014 at 8km. What are the plasma frequencies
there?
1
Expert's answer
2020-07-01T15:42:19-0400

The plasma frequencies can be expressed as


"f=\frac{\omega}{2\pi}=\frac{e}{2\pi}\sqrt{\frac{n_e}{m_e\epsilon_0}}."

At 50-km altitude: "f=8.979\text{ GHz}"

At 70 km: "f=2.839\text{ GHz}"

At 8 km: "f=89.79\text{ MHz}"